OpenStudy (anonymous):

Quick question about graphing calculator. Writing below:

OpenStudy (anonymous):

I have a TI-83 Plus. Here's the question: On the "Y=" screen you have Y1, Y2, Y3, etc. If you have multiple graphs entered, how can you choose just one of the graph to show up on the graphing screen?

OpenStudy (anonymous):

Remove 'delete' the other two equations from the "Y=" screen.

OpenStudy (anonymous):

well i am doing composite functions so i have an equation for y1, a differnt equation for y2, and then my y3 is "y=(y1)+(y2)

OpenStudy (anonymous):

i just want to show the graph of "y=(y1)+(y2)"

OpenStudy (anonymous):

That I'm not sure how to do that, I have a Ti-84+. Is there a reason why you want just the "y=(y1)+(y2)"?

OpenStudy (anonymous):

because we are doing composite functions so i just want the composite graph. For example: f(x)=2x+3 g(x)=(x-3)^2 find f(g(x))

OpenStudy (anonymous):

so i want to check my answers by using the calc. see what i mean?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

Ok I think there is an option you can do, if you hit trace and then use the up/down arrow keys you can switch which equation you want.

OpenStudy (anonymous):

hmmm i see what you are saying but i know there is a way to have just one graph appear

OpenStudy (anonymous):

Well sorry, you could take a look on Texas Instruments website to find an answer or your calculator manual.

OpenStudy (anonymous):

thanks for your help, ill go check there

hero (hero):

You can chat with a Texas Instruments rep or just call them I'm pretty sure they will speak to you for free about it.